OpenStudy (anonymous):

Find an exact value of sin(19pi.2)

OpenStudy (anonymous):

\[\sin(\frac{19\pi}{2})=\sin(\frac{16\pi}{2}+\frac{3\pi}{2})=\sin(8\pi+\frac{3\pi}{2})=\sin(\frac{3\pi}{2})\]